OpenStudy (anonymous):

Calculate the average rate of change for the given function, from x = 2 to x = 7.

OpenStudy (kirbykirby):

\[\text{rate of change}=\frac{\Delta y}{\Delta x}=\frac{\Delta f(x)}{\Delta x} =\frac{f(x_2)-f(x_1)}{x_2-x_1}\]

OpenStudy (kirbykirby):

\(x_1=2\), \(x_2=7\)

OpenStudy (anonymous):

thank